# Pedestrian Crossing Discrepancy Within Static and Dynamic Crowds: An Experimental Study

### v1

@@ 4.2.1 Crossing Efficiency
The average crossing speed ($\bar{v}_{x}$ ) denotes the average velocity in the crossing direction within the crossing process, and serves as the measure of pedestrians’ cross-efficiency within the crowd, expressed as:
@@
Here, $L$ denotes the crossing length, $t_{{\rm{in}}}$ indicates the moment of the pedestrian entering the experimental area, and $t_{{\rm{out}}}$ represents the moment of the pedestrian leaving the experimental area. The distribution of average crossing speeds under static and dynamic configurations is presented in Fig.9. It is evident that the average crossing speed of the static crowd is significantly higher than the dynamic crowd. Within the static condition, the highest speed peak is observed in experiment S-L5-N5 (mean = 1.59 m/s), whereas the dynamic condition exhibits its speed peak in experiment D-L5-N0 (mean = 1.48 m/s). As the global density increases, the disparity in average crossing speed tends to widen.
Figure 9: Pedestrians’ average crossing speed variation in different experimental groups, with median (thick green line and green numbers), upper and lower quartiles (yellow dotted line), and mean data (red and blue numbers). The subplots reflected the trend of the speed difference in static and dynamic contexts and the variation trend of mean and standard deviation(mean±SD), respectively.
@@ 4.2.2 Orthogonal Analysis
In this part, the pedestrians’ velocity is decomposed into two orthogonal components: the crossing direction (in the $x$-direction) and the transverse direction (in the $y$-direction). This decomposition allows us to investigate the variations in crossing velocity ($v_{x}$) and transverse velocity ($v_{y}$) so that different detour features within static and dynamic contexts can be quantitatively elucidated. The velocity deconstruction correspondence within each experiment is independently analyzed, as illustrated in Fig.10.
Figure 10: The kernel density and the associated marginal histograms of instantaneous velocities along orthogonal directions for various experiments.
@@
[…]
Fig. 11 illustrates the corresponding variations in the instantaneous velocity ($v$), the crossing velocity ($v_{x}$), and the absolute value of the transverse velocity ($|v_{y}|$). A comparison between Fig.12(a) and Fig.12(b) reveals that velocity fluctuations expand with increasing global density. Moreover, within the dynamic context, the fluctuations of the crossing velocity are more pronounced compared to the static context. Most importantly, in both dynamic and static experiments, the dynamics of pedestrians crossing exhibit a fundamental pattern: with an increase in global density, the instantaneous velocity and crossing velocity decrease, while the transverse velocity remains within the interval of [0.2,0.5]. This observation implies that even in highly congested scenarios, pedestrians will continue to strategically keep motion through detours (Wang et al., 2023).
Figure 12: Relationship between instantaneous velocity and component velocity in different experimental groups. The dash and the shaded area represent the mean value and one-sigma error (mean ± standard deviation), respectively.
@@ 5 Conclusions
In this paper, we investigated pedestrian crossing behavior within dynamic and static contexts. The qualitative analysis of pedestrian trajectories reveals distinct characteristics when pedestrians crossing static or dynamic crowds. The formation of cross-channels, a unique self-organization phenomenon within static crowds, has been observed.
To investigate the discrepancies in pedestrian crossing behavior, we introduced quantitative metrics to analyze the differences in motion patterns and trajectories. Within static contexts, pedestrians’ speeds and swarm factors are not sensitive to the increases in global density and do not exhibit typical behavioral patterns. The similar desensitization mechanism might contribute to the formation of cross-channels. In contrast, within dynamic crowds, the speeds and swarm factors of crossing pedestrians show dramatic fluctuations and contribute to quantitatively classifying the anticipation and reaction behaviors. Moreover, the orthogonal analysis of velocity illustrated the cross-efficiency variation induced by the crowd contexts. Fundamental patterns in pedestrian motion have been identified: with the elevation of global density, both the instantaneous and crossing velocities decrease, while transverse velocity appears unaffected. This phenomenon proves detour behavior may serve as a crucial mechanism for pedestrians to maintain dynamism within dense crowds.
This study contributes to the empirical knowledge of pedestrian behavior pattern and a deeper understanding of crowd dynamics.

### v2

@@ 4.2.1 Crossing Efficiency
The average crossing speed ($\bar{v}_{x}$) denotes the average speed in the crossing direction and serves as the measure of crossing efficiency within the crowd, expressed as:
${\bar{v}_{x}}=\frac{L}{{{t_{{\rm{out}}}}-{t_{{\rm{in}}}}}}.$ (3)
Here, $L$ denotes the crossing length, $t_{{\rm{in}}}$ indicates the moment of the pedestrian entering the experimental area, and $t_{{\rm{out}}}$ represents the moment of the pedestrian leaving the experimental area. The distribution of average crossing speeds under static and dynamic contexts is presented in Fig.7. The average crossing speed of the static crowd is significantly higher than the dynamic crowd. Within the static condition, the highest speed peak is observed in experiment S-L5-N5 (mean = 1.59 m/s), whereas the dynamic condition exhibits its speed peak in experiment D-L5-N0 (mean = 1.48 m/s). As global density increases (within 0-2 ped/m2), the disparity in average crossing speed tends to widen, as shown in the subplots.
@@ 4.2.2 Orthogonal Analysis
Figure 8: Scatter plots and the corresponding marginal histograms of instantaneous velocities along orthogonal directions, *and ** indicate that, according to the Bartlett test, the transverse velocity data from static and dynamic contexts have significantly different variances at the 0.05 level and 0.01 level, respectively.
In this section, the instantaneous velocity is decomposed into two orthogonal components: the crossing direction (in the $x$-direction) and the transverse direction (in the $y$-direction). This decomposition allows us to investigate the variations in crossing velocity ($v_{x}$) and transverse velocity ($v_{y}$) so that different detour features within static and dynamic contexts can be quantitatively elucidated. The velocity deconstruction correspondence within each experiment is independently analyzed, as illustrated in Fig.8. The results of Bartlett’s test for homogeneity of variances in transverse velocity across experiments are presented in Tab.2. Fig.8 shows a consistent linear relationship between the crossing velocity ($v_{x}$) and the instantaneous velocity ($v$). Moreover, by comparing the distribution trends at different global densities, the transverse velocity ($v_{y}$) exhibits a systematic increase within the dynamic context (also see Tab.2), indicating that pedestrians have a higher frequency of detours.
@@
Table 2: Results of Bartlett’s test for homogeneity of variances, NS: not significant.
Index $\chi^{2}$ $p$-value Index $\chi^{2}$ $p$-value Index $\chi^{2}$ $p$-value
S(D)-L5-N5 25.6737 $p<0.01$ S(D)-L5-N30 0.5345 NS S(D)-L3-N18 0.1575 NS
S(D)-L5-N10 22.1530 $p<0.01$ S(D)-L5-N35 21.6510 $p<0.01$ S(D)-L3-N24 70.1284 $p<0.01$
S(D)-L5-N15 0.7035 NS S(D)-L5-N40 3.7621 NS S(D)-L3-N30 4.3785 $p<0.05$
S(D)-L5-N20 30.1186 $p<0.01$ S(D)-L5-N45 16.3658 $p<0.01$ S(D)-L3-N36 6.4991 $p<0.05$
S(D)-L5-N25 5.1891 $p<0.05$ S(D)-L5-N49 1.0561 NS S(D)-L3-N42 0.0793 NS
Fig.10 illustrates the corresponding variations in the instantaneous velocity ($v$), the crossing velocity ($v_{x}$), and the absolute value of transverse velocity ($|v_{y}|$). A comparison between Fig.10(a) and Fig.10(b) reveals that speed fluctuations expand with increasing global density. Moreover, in dynamic contexts, the decay trend of the crossing velocity (including instantaneous velocity) is more pronounced than in static contexts, as confirmed by statistical analysis. Most importantly, in both dynamic and static experiments, variations in speeds exhibit a fundamental pattern: as global density increases, both instantaneous velocity and crossing velocity decrease, while transverse velocity primarily remains within the interval of [0.2, 0.5]. This observation implies that even in highly congested scenarios, pedestrians may continue to strategically maintain motion through detours (Wang et al., 2023).
Figure 10: Variation of instantaneous velocity ($v$), crossing velocity ($v_{x}$) and the absolute value of transverse velocity $|v_{y}|$ versus global density. The dash and the shaded area represent the mean value and one-sigma error, respectively.
@@ 5 Conclusions
This study investigated pedestrian crossing behavior within dynamic and static crowds. To explore discrepancies in pedestrian crossing behavior, a quantitative density metric termed the swarm factor was introduced. In static contexts, the limited variation in speed and swarm factor suggests that crossing pedestrians are less responsive to density increases, indicating potential tactical behavior that may contributes to the formation of cross-channels, a unique self-organization phenomenon. This contrasts with dynamic crowds, where speed and swarm factor synchronization inversely, indicating density-dependent behavioral adaptation (operational level). Furthermore, employing orthogonal analysis, we identified a fundamental pattern in pedestrian crossing: as global density increases, both instantaneous velocity and crossing velocity decrease, while transverse velocity remains stable. This finding suggests that detour behavior may be a potential mechanism for pedestrians to maintain dynamism within crowds. These observations contribute to the knowledge of pedestrian behavior patterns and provide an empirical understanding of crowd dynamics.